An investment will pay $100 at the end of each of the next 3 years, $250 at the end of Year 4, $350 at the end of Year 5, and $500 at the end of Year 6. If other investments of equal risk earn 11% annually, what is its present value? Its future value? Do not round intermediate calculations. Round your answers to the nearest cent.
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=100/1.11+100/1.11^2+100/1.11^3+250/1.11^4+350/1.11^5+500/1.11^6

=$884.08(Approx).

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value=884.08*(1.11)^6

=$1653.6(Approx).
Answer #2

SOLUTION :


Discount rate, r = 11% = 0.11 

=> 1 + r = 1.11


PV 

= 100/1.11+100/1.11^2+100/1.11^3+250/1.11^4+350/1.11^5+500/1.11^6

= 884.08 ($) (ANSWER).


FV 

= PV*1.06^6 

= 884.08*1.11^6

= 1653.60 ($) (ANSWER)

answered by: Tulsiram Garg

> Under FV, first line, please correct " = PV*1.06^6 " as " = PV * 1.11^6 "

Tulsiram Garg Thu, Oct 14, 2021 3:09 AM